OpenStudy (anonymous):

A recipe calls for 2 tsp of butter. Melinda wants to make 40% more servings than the recipe makes. Which expression represents how many teaspoons of butter Melinda should use? A. 0.6 * 2 B. 1.6 * 2 C. 0.4 * 2 D. 1.4 * 2

OpenStudy (anonymous):

Ok what questions do you know that are not true

OpenStudy (anonymous):

c and d

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

ok good

OpenStudy (michele_laino):

I think that your answer is D. because we have to compute this: 2+(0.4*2)=2+0.8=2.8=2*1.4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i thought it was b

OpenStudy (anonymous):

@Secret-Ninja

OpenStudy (secret-ninja):

I guessed C but I don't know, I keep having brain farts today, I can't think straight. >.< I don't recommend asking me today lol

OpenStudy (anonymous):

kk XD